A ball bearingX is dropped vertically downwards from the edge of a table and it takes 0.5 seconds to hit the floor.another one y leaves the edge of the table horizontally with a velocity of 5mpr second find
A.time taken for y to reach the floor
1
Expert's answer
2020-04-02T11:01:06-0400

As per the given question,

time taken to reach the ground by X one "(t_1)=0.5 sec"

we know that "h=ut+\dfrac{gt^2}{2}"

u=0

"h=0+\dfrac{9.8\times 0.5 \times 0.5}{2}"

"h=1.225m"

now for the Y,

object is projected in horizontal direction, but there is no vertical velocity, so intital velocity in the vertical direction =0

so "h=ut+\dfrac{gt^2}{2}"

"\Rightarrow 1.225=0+\dfrac{9.8\times t^2}{2}"

"t=\sqrt{0.25}=0.5 sec"